OpenStudy (anonymous):

reference angle for -13pi/5?

OpenStudy (lasttccasey):

Well if you're looking for the angle on the reference circle that relates to -13 pi/4, just keep adding 2 until the number is between 0 and 2pi. -13/4 + 2 = - 1.25 -1.25 + 2 = 0.75 3pi/4 should be your reference angle for a standard circle with 0 on the right horizontal.

OpenStudy (jdoe0001):

do you know what a "negative angle" is anyway?

OpenStudy (lasttccasey):

It's just an angle going clockwise instead of the typical counterclockwise. you can go 180 (pi) or -180 (-pi) degrees and you'll be at the same place on a circle.

OpenStudy (anonymous):

thanks a lot for your help!!!!